OpenStudy (theraggedydoctor):

A car traveled 200 miles at an average speed of 45 miles per hour. Of the following, which is the closest approximation to the amount of time that could be saved on this 200-mile trip if the average speed had increased 20 percent? 1 hour 3/4 hour 1/2 hour 1/4 hour 1/5 hour

OpenStudy (anonymous):

s=vt 200=45t t=200/45=40/9 hours new speed v=45*(1+(20/100))=45*(1+(1/5))=45*6/5=54 200=54t t=200/54=100/27 hours\[difference~ intime=\frac{ 40 }{ 9 }-\frac{ 100 }{ 27 }=\frac{ 120-100 }{2 7 }=\frac{ 20 }{ 27 }\] \[=\frac{ 3 }{ 4 } ~ approximately\]

OpenStudy (theraggedydoctor):

Thank you! c:

OpenStudy (anonymous):

yw